Concert tickets cost $2.00/child and $5.00/adult. A group of thirty people
consisting of children and adults paid a total of $87.00 for the concert. How
many adults were in the group?
: Concert tickets cost $2.00/child and $5.00/adult. A group of thirty people
consisting of children and adults paid a total of $87.00 for the concert. How
many adults were in the group?

Answer by Fombitz(1740)

C-Child
A-Adult
1.A+C=30Total tickets sold.
2.5A+2C=87 Total tickets cost.
Use eq. 1 to solve for A as a function of C.
1.A+C=30
A=30-C
Substitute this expression for A in eq. 2,
2.5A+2C=87
5(30-C)+2C=87
Solve for C.
150-5C+2C=87
-3C=-63
C=21
Use eq. 1 to solve for A.
A=30-C
A=30-21
A=9
9 adult and 21 children tickets were sold.